# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 1 Introduction
Compressed sensing has proven to be an effective method in acquiring and reconstructing high-dimensional signals [20, 17, 6, 5, 13]. Mathematically, the goal of linear compressed sensing is to reconstruct sparse signals $\mathbf{x}$ from a set of measurements $\mathbf{y}=\mathbf{A}\mathbf{x}+\bm{\epsilon}$, under the sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ and noise vector $\bm{\epsilon}\in\mathbb{R}^{m}$. Restricted isometry property (RIP) lies at the center of linear compressed sensing theory, whose major finding is a set of efficient algorithms achieving instance optimality under RIP sensing matrices [20, 47, 51, 2, 14, 5]. Here, the instance optimality describes the capacity of an algorithm to achieve estimation error proportional to the signal’s distance to the cone of $s$-sparse vectors $\Sigma^{n}_{s}$. In the noiseless case, this translates into exact reconstruction of sparse signals and accurate estimate of approximately sparse signals. As an example, if $\mathbf{A}$ satisfies RIP over the cone of sparse vectors and $\varepsilon\geq\|\bm{\epsilon}\|_{2}$, then basis pursuit $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$
@@
for some absolute constants $C_{1},C_{2}$, where $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s}):=\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{x}-\mathbf{u}\|_{1}$ denotes the $\ell_{1}$ distance of $\mathbf{x}$ to $\Sigma^{n}_{s}.$ More details are given in Section 2.
The focus of the present paper is on the nonlinear compressed sensing model of phase-only compressed sensing (PO-CS), which concerns the reconstruction of sparse signals in the standard sphere $\mathbb{S}^{n-1}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|_{2}=1\}$ from
@@
under a complex sensing matrix $\bm{\Phi}=[\bm{\Phi}_{1},\cdots,\bm{\Phi}_{m}]^{*}\in\mathbb{C}^{m\times n}$, with the phase function being given by $\operatorname{sign}(c)=\frac{c}{|c|}$ for $c\in\mathbb{C}\setminus\{0\}$ and $\operatorname{sign}(0)=1$ by convention. We assume that we observe $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ that equaling $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ in the noiseless case or a perturbed version of $\mathbf{z}$ in the noisy case. We note that phase-only reconstruction of unstructured signals has been well studied [26, 38, 9, 25, 18, 32, 48, 33].
PO-CS was initially considered as a natural extension of 1-bit compressed sensing [^1] This concerns the recovery of $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})$ with real sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}.$ [3, 4] and was recently revisited in [19, 10, 27, 8]. Exact reconstruction was observed experimentally in [4] and theoretically proved very recently by Jacques and Feuillen [27] who proposed to recast PO-CS as a linear compressed sensing problem. We proceed to introduce this linearization approach under the noiseless setting $\breve{\mathbf{z}}=\mathbf{z}$. Since $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ implies that the entries of $\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi}\mathbf{x}$ are non-negative real numbers, the phases give the linear measurements
@@
[…]
@@
In [10], the authors used a covering argument to show that the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}:=\Sigma^{n}_{s}\cap\mathbb{S}^{n-1}\}$ simultaneously obey RIP under a near-optimal number of measurements. This leads to a uniform exact reconstruction guarantee over $\Sigma^{n,*}_{s}$, but provides no guarantee for $\mathbf{x}\notin\Sigma^{n,*}_{s}$. Thus, their guarantee is not instance optimal.
@@
The main ingredient is to show the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$, where $\mathbb{B}_{1}^{n}(\sqrt{2s})=\{\mathbf{u}\in\mathbb{R}^{n}:\|\mathbf{u}\|_{1}\leq\sqrt{2s}\}$ is the scaled-$\ell_{1}$ ball, simultaneously satisfy RIP. Note that [10] proved that the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy RIP through a covering argument, yet their arguments are not sufficient to prove the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$. The main issue is that $\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}$ is essentially larger than $\Sigma^{n,*}_{s}$ in terms of metric entropy (or covering number) under an $o(1)$ covering radius. To that end, we utilize a finer treatment to the perturbation of the complex phases to avoid a heavy-tailed random process. See Appendix A.1 for a summary. More generally, we establish the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ over some cone $\mathcal{U}$ for an arbitrary set $\mathcal{K}\subset\mathbb{S}^{n-1}$. Here is an informal version.
@@
Given a cone $\mathcal{U}$ in $\mathbb{R}^{n}$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if $m\geq C_{1}\big{(}\omega^{2}(\mathcal{U}\cap\mathbb{S}^{n-1})+\omega^{2}(\mathcal{K})\big{)}$ with sufficiently large $C_{1}$, then with high probability on the complex Gaussian $\bm{\Phi}$, the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ satisfy RIP over $\mathcal{U}$ with small enough distortion.
@@
Our results provide answers to all of these questions. Let us consider the nonuniform recovery of a fixed sparse signal. We show that small dense noise $\bm{\tau}$ considered in [10, 27], even when appearing before taking the phases (i.e., pre-sign noise), increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\tau_{0})$. Moreover, the $O(\tau_{0})$ bound achieved by $\mathbf{x}^{\sharp}$ for pre-sign/post-sign dense noise is nearly tight over all algorithms. We also investigate the impact of sparse corruption which adversarially moves $\zeta_{0}m$ measurements to arbitrary phase-only values. This increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ if $\zeta_{0}$ is small enough. We expect that $\tilde{O}(\sqrt{\zeta_{0}})$ is tight for the specific estimator $\mathbf{x}^{\sharp}$, which we support by providing a partial analysis and numerical evidence. However, for general estimators such dependence is suboptimal and can be improved to “zero,” as in this regime we can still exactly recover $\mathbf{x}$, akin to existing studies of corrupted linear systems and corrupted sensing [21, 7, 24]. In particular, we propose to reformulate PO-CS under sparse corruption to a noiseless linear compressed sensing problem with an extended new sensing matrix. This matrix is then shown to satisfy RIP, thus implying exact reconstruction. See Table 1 for a summary of these results.
@@
Sparse Corruption $\bm{\zeta}$ $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ $O\Big{(}\sqrt{\zeta_{0}\log(1/\zeta_{0})}\Big{)}$ Thm. 3.6 Partially Yes Prop. 3.1 No, 00 is optimal Thm. 3.8 Fig. 3
@@
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\tau_{0}+C_{3}\sqrt{\zeta_{0}\log(1/\zeta_{0})}+C_{4}\sqrt{\frac{s\log(\frac{en}{s})}{m}},\qquad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (1.9)
@@ 2 Preliminaries
We start with some notational conventions. We denote matrices and vectors by boldface letters, and scalars by regular letters. $|\mathcal{S}|$ denotes the cardinality of a finite set $\mathcal{S}$. We use $\log(\cdot)$ to denote the natural logarithm to the base of the mathematical constant $e$. The standard Euclidean sphere, the $\ell_{2}$-ball and the $\ell_{1}$-ball in $\mathbb{R}^{n}$ are denoted by $\mathbb{S}^{n-1}$, $\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}$, respectively. Given $\mathcal{K},\mathcal{K}^{\prime}\subset\mathbb{R}^{n}$ and some $\lambda\in\mathbb{R}$, we let $\mathcal{K}+\lambda\mathcal{K}^{\prime}:=\{\mathbf{u}+\lambda\mathbf{v}:\mathbf{u}\in\mathcal{K},\mathbf{v}\in\mathcal{K}^{\prime}\}$, $\mathcal{K}_{(\lambda)}=(\mathcal{K}-\mathcal{K})\cap(\lambda\mathbb{B}_{2}^{n})$ and $\mathcal{K}^{\rm(N)}=\mathcal{K}\cap\mathbb{S}^{n-1}$. We write $\mathbb{B}_{2}^{n}(\mathbf{u};r):=\mathbf{u}+r\mathbb{B}_{2}^{n}$, $\mathbb{B}_{2}^{n}(r):=r\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}(r):=r\mathbb{B}_{1}^{n}$. We also define $\operatorname{rad}(\mathcal{K})=\sup_{\mathbf{u}\in\mathcal{K}}\|\mathbf{u}\|_{2}$. Recall that $\Sigma^{n,*}_{s}:=(\Sigma^{n}_{s})^{\rm(N)}$ is the set of all $s$-sparse signals in $\mathbb{S}^{n-1}$.
We refer to complex numbers with absolute value $1$ as the phase-only values. For a vector $\mathbf{u}=[u_{i}]\in\mathbb{C}^{n}$, we work with the $\ell_{p}$-norm $\|\mathbf{u}\|_{p}=(\sum_{i}|u_{i}|^{p})^{1/p}$ ($p\geq 1$), max norm $\|\mathbf{u}\|_{\infty}=\max_{i}|u_{i}|$, and zero “norm” $\|\mathbf{u}\|_{0}$ that counts the number of non-zero entries. Further given $\mathbf{v}=[v_{i}]\in\mathbb{C}^{n}$, we work with the inner product $\langle\mathbf{u},\mathbf{v}\rangle=\mathbf{u}^{*}\mathbf{v}=\sum_{i=1}^{n}\overline{u_{i}}v_{i}$ and the Hadamard product $\mathbf{u}\odot\mathbf{v}=(u_{1}v_{1},u_{2}v_{2},\cdots,u_{n}v_{n})^{\top}$. For a complex matrix $\mathbf{A}=\mathbf{B}+\mathbf{C}\mathbf{i}$ with $\mathbf{i}$ reserved for the complex unit, we will use $\mathbf{A}^{\Re}$ or $\Re(\mathbf{A})$ to denote its real part $\mathbf{B}$, and $\mathbf{A}^{\Im}$ or $\Im(\mathbf{A})$ to denote its imaginary part $\mathbf{C}$. For a random variable $X$ we define the sub-Gaussian norm as $\|X\|_{\psi_{2}}=\inf\{t>0:\mathbbm{E}\exp(X^{2}/t^{2})<2\}$ and the sub-exponential norm as $\|X\|_{\psi_{1}}=\inf\{t>0:\mathbbm{E}\exp(|X|/t)<2\}$. For independent zero-mean sub-Gaussian variables $\{X_{i}\}_{i=1}^{N}$, there exists absolute constant $C$ such that
@@
The sub-Gaussian norm of a random vector $\mathbf{X}\in\mathbb{R}^{n}$ is defined as $\|\mathbf{X}\|_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{S}^{n-1}}\|\mathbf{v}^{\top}\mathbf{X}\|_{\psi_{2}}$. We refer readers to [49, Sec. 2] for details on these definitions and properties.
We use $\{C,C_{1},C_{2},\cdots\}$ and $\{c,c_{1},c_{2},\cdots\}$ to denote absolute constants whose values may vary from line to line. For two positive quantities $I_{1}$ and $I_{2}$, we write $I_{1}=O(I_{2})$ if $I_{1}\leq CI_{2}$ holds for some $C$, and write $I_{1}=\Omega(I_{2})$ if $I_{1}\geq cI_{2}$ for some $c>0$. We write $I_{1}=\Theta(I_{2})$ if $I_{1}=O(I_{2})$ and $I_{1}=\Omega(I_{2})$ simultaneously hold. We also use $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),\tilde{\Theta}(\cdot)$ as the less precise versions of these that hide log factors. We use $o(1)$ to generically denote quantity that tends to zero when $m,n,s\to\infty$.
We say $I_{1}$ is small enough (or sufficiently small) if $I_{1}\leq c_{1}$ for some suitably small constant $c_{1}$. Conversely, it is large enough (or sufficiently large) if $I_{1}\geq C_{1}$ for some suitably large constant $C_{1}$. We say $I_{1}$ is bounded away from 00 if $I_{1}\geq c_{1}$ for some $c_{1}>0$.
Linear Compressed Sensing: Let $\mathcal{U}$ be a cone in $\mathbb{R}^{n}$, we say $\mathbf{A}$ satisfies RIP over $\mathcal{U}$ with distortion $\delta>0$, denoted by $\mathbf{A}\sim\operatorname{RIP}(\mathcal{U},\delta)$, if

### v3

@@ 1 Introduction
Compressed sensing has proven to be an effective method in acquiring and reconstructing high-dimensional signals [20, 17, 6, 5, 13]. Mathematically, the goal of linear compressed sensing is to reconstruct sparse signals $\mathbf{x}$ from a set of measurements $\mathbf{y}=\mathbf{A}\mathbf{x}+\bm{\epsilon}$, under the sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ and noise vector $\bm{\epsilon}\in\mathbb{R}^{m}$. Restricted isometry property (RIP) lies at the center of linear compressed sensing theory, whose major finding is a set of efficient algorithms achieving instance optimality under RIP sensing matrices [20, 48, 52, 2, 14, 5]. Here, the instance optimality describes the capacity of an algorithm to achieve estimation error proportional to the signal’s distance to the cone of $s$-sparse vectors $\Sigma^{n}_{s}$. In the noiseless case, this translates into exact reconstruction of sparse signals and accurate estimate of approximately sparse signals. As an example, if $\mathbf{A}$ satisfies RIP over the cone of sparse vectors and $\varepsilon\geq\|\bm{\epsilon}\|_{2}$, then basis pursuit
@@
for some absolute constants $C_{1},C_{2}$, where $\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s}):=\min_{\mathbf{u}\in\Sigma^{n}_{s}}\|\mathbf{x}-\mathbf{u}\|_{1}$ denotes the $\ell_{1}$ distance of $\mathbf{x}$ to $\Sigma^{n}_{s}.$ More details are given in Section 2. We write $\Delta(\mathbf{A};\mathbf{y};\varepsilon)$ for the estimator $\hat{\mathbf{x}}$ in (1.1) to emphasize the required inputs $(\mathbf{A},\mathbf{y},\varepsilon).$
The focus of the present paper is on the nonlinear compressed sensing model of phase-only compressed sensing (PO-CS), which concerns the reconstruction of sparse signals in the standard sphere $\mathbb{S}^{n-1}=\{\mathbf{x}\in\mathbb{R}^{n}:\|\mathbf{x}\|_{2}=1\}$ from [^1] We focus on real-valued signals in this paper, but point out that our results generalize to complex-valued signals by slightly more work (see [10]).
@@
under a complex sensing matrix $\bm{\Phi}=[\bm{\Phi}_{1},\cdots,\bm{\Phi}_{m}]^{*}\in\mathbb{C}^{m\times n}$, with the phase function being given by $\operatorname{sign}(c)=\frac{c}{|c|}$ for $c\in\mathbb{C}\setminus\{0\}$ and $\operatorname{sign}(0)=1$ by convention. Let $\breve{\mathbf{z}}\in\mathbb{C}^{m}$ be our observation vector, which equals $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ in the noiseless case and is a perturbed version of $\mathbf{z}$ in the noisy case. [^2] Throughout this paper, $\mathbf{x}$ denotes the underlying true signal, $\bm{\Phi}\in\mathbb{C}^{m\times n}$ denotes the complex sensing matrix, $\mathbf{z}$ is reserved for the noiseless phase-only observations $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$, and $\breve{\mathbf{z}}$ is our observations and a possibly noisy version of $\mathbf{z}$. Therefore, $\mathbf{z}$ and $\breve{\mathbf{z}}$ implicitly depend on $\mathbf{x}$. Such a model may find applications in sensing systems subject to severe dynamic-range limitations or unknown gain fluctuations, and may also be useful under extreme quantization or multiplicative noise, since phase lies on a compact unit circle and is invariant under unknown scaling. We note that phase-only reconstruction of unstructured signals was studied extensively several decades ago [25, 39, 9, 26, 18, 33, 49, 34].
PO-CS was initially considered as a natural extension of 1-bit compressed sensing [^3] This concerns the recovery of $\mathbf{x}\in\Sigma^{n,*}_{s}$ from $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})$ with real sensing matrix $\mathbf{A}\in\mathbb{R}^{m\times n}.$ Note that the phase-only measurements (1.3) take continuous values (on the complex unit circle) and are therefore more informative than 1-bit measurements $\mathbf{y}=\operatorname{sign}(\mathbf{A}\mathbf{x})\in\{-1,1\}^{m}$. This distinction leads to a fundamental difference between the two problems: while 1-bit compressed sensing admits a strictly positive information-theoretic lower bound on the reconstruction error [28], exact recovery can be achieved in phase-only compressed sensing. [3, 4] and was recently revisited in [19, 10, 27, 8]. Exact reconstruction was observed experimentally in [4] and theoretically proved very recently by Jacques and Feuillen [27] who proposed to recast PO-CS as a linear compressed sensing problem. We proceed to introduce this linearization approach under the noiseless setting $\breve{\mathbf{z}}=\mathbf{z}$. Since $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$ implies that the entries of $\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi}\mathbf{x}$ are non-negative real numbers, the phases give the linear measurements
@@
$\text{find sparse}~\mathbf{u},~~\text{such that }\begin{bmatrix}\frac{1}{\kappa m}\Re(\breve{\mathbf{z}}^{*}\bm{\Phi})\\ \frac{1}{\sqrt{m}}\Im(\operatorname{diag}(\breve{\mathbf{z}}^{*})\bm{\Phi})\end{bmatrix}\mathbf{u}=\mathbf{e}_{1}.$ (1.6)
@@
In [10], the authors used a covering argument to show that the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\Sigma^{n,*}_{s}:=\Sigma^{n}_{s}\cap\mathbb{S}^{n-1}\}$ simultaneously obey RIP under a near-optimal number of measurements. This leads to a uniform exact reconstruction guarantee over $\Sigma^{n,*}_{s},$ but provides no guarantee for $\mathbf{x}\notin\Sigma^{n,*}_{s}$. Thus, their guarantee is not instance optimal.
@@
The main ingredient is to show the new sensing matrices in $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$, where $\mathbb{B}_{1}^{n}(\sqrt{2s})=\{\mathbf{u}\in\mathbb{R}^{n}:\|\mathbf{u}\|_{1}\leq\sqrt{2s}\}$ is the scaled-$\ell_{1}$ ball, simultaneously satisfy RIP. Note that [10] proved that the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\Sigma^{n,*}_{s}\}$ simultaneously satisfy RIP through a covering argument, yet their arguments are not sufficient to prove the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x}),\,\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}\}$. The main issue is that $\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}$ is essentially larger than $\Sigma^{n,*}_{s}$ in terms of metric entropy (or covering number) under an $o(1)$ covering radius. To that end, we utilize a finer treatment to the perturbation of the complex phases to avoid a heavy-tailed random process. See Appendix A.1 for a summary. More generally, we establish the RIP of $\{\mathbf{A}_{\mathbf{z}}:\mathbf{x}\in\mathcal{K}\}$ over some cone $\mathcal{U}$ for an arbitrary set $\mathcal{K}\subset\mathbb{S}^{n-1}$. Here is an informal version, where $\omega(\mathcal{U})$ denotes the Gaussian width of a set $\mathcal{U}$ (see Section 2.2).
@@
Given a cone $\mathcal{U}$ in $\mathbb{R}^{n}$ and $\mathcal{K}\subset\mathbb{S}^{n-1}$, if
$m\geq C_{1}\big(\omega^{2}(\mathcal{U}\cap\mathbb{S}^{n-1})+\omega^{2}(\mathcal{K})\big)$
with sufficiently large $C_{1}$, then with high probability on the complex Gaussian $\bm{\Phi}$, the matrices $\{\mathbf{A}_{\mathbf{z}}:\mathbf{z}=\operatorname{sign}(\mathbf{A}\mathbf{x}),\,\mathbf{x}\in\mathcal{K}\}$ satisfy RIP over $\mathcal{U}$ with small enough distortion.
@@
Our results provide answers to all of these questions. Let us consider the nonuniform recovery of a fixed sparse signal. We show that small dense noise $\bm{\tau}$ considered in [10, 27], even when appearing before taking the phases (i.e., pre-sign noise), increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\tau_{0})$. Moreover, the $O(\tau_{0})$ bound achieved by $\mathbf{x}^{\sharp}$ for pre-sign/post-sign dense noise is nearly tight over all algorithms. We also investigate the impact of sparse corruption which adversarially moves $\zeta_{0}m$ measurements to arbitrary phase-only values. This increments $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}$ by $O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ if $\zeta_{0}$ is small enough. We expect that $\tilde{O}(\sqrt{\zeta_{0}})$ is tight for the specific estimator $\mathbf{x}^{\sharp}$, which we support by providing a partial analysis and numerical evidence. However, for general estimators such dependence is suboptimal and can be improved to “zero,” as in this regime we can still exactly recover $\mathbf{x}$. In particular, we propose to reformulate PO-CS under sparse corruption to a noiseless linear compressed sensing problem with an extended new sensing matrix. This matrix is then shown to satisfy RIP, thus implying exact reconstruction. See Table 1 for a summary of these results.
@@
Sparse Corruption $\bm{\zeta}$ $\|\bm{\zeta}\|_{0}\leq\zeta_{0}m$ $O\Big(\sqrt{\zeta_{0}\log(1/\zeta_{0})}\Big)$ Thm. 3.6 Partially Yes Prop. 3.1 No, 0 is optimal Thm. 3.8 Fig. 3
@@
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq C_{1}\frac{\sigma_{\ell_{1}}(\mathbf{x},\Sigma^{n}_{s})}{\sqrt{s}}+C_{2}\tau_{0}+C_{3}\sqrt{\zeta_{0}\log(1/\zeta_{0})}+C_{4}\sqrt{\frac{s\log(\frac{en}{s})}{m}}\mathbbm{1}(\zeta_{0}>0),\quad\forall\mathbf{x}\in\mathbb{S}^{n-1}.$ (1.9)
@@ 2 Preliminaries
We start with some notational conventions. We denote matrices and vectors by boldface letters, and scalars by regular letters. $|\mathcal{S}|$ denotes the cardinality of a finite set $\mathcal{S}$. We use $\log(\cdot)$ to denote the natural logarithm to the base of the mathematical constant $e$. The standard Euclidean sphere, the $\ell_{2}$-ball and the $\ell_{1}$-ball in $\mathbb{R}^{n}$ are denoted by $\mathbb{S}^{n-1}$, $\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}$, respectively. Given $\mathcal{K},\mathcal{K}^{\prime}\subset\mathbb{R}^{n}$ and some $\lambda\in\mathbb{R}$, we let $\mathcal{K}+\lambda\mathcal{K}^{\prime}:=\{\mathbf{u}+\lambda\mathbf{v}:\mathbf{u}\in\mathcal{K},\mathbf{v}\in\mathcal{K}^{\prime}\}$, $\mathcal{K}_{(\lambda)}=(\mathcal{K}-\mathcal{K})\cap(\lambda\mathbb{B}_{2}^{n})$ and $\mathcal{K}^{\mathbb{S}}=\mathcal{K}\cap\mathbb{S}^{n-1}$. We write $\mathbb{B}_{2}^{n}(\mathbf{u};r):=\mathbf{u}+r\mathbb{B}_{2}^{n}$, $\mathbb{B}_{2}^{n}(r):=r\mathbb{B}_{2}^{n}$ and $\mathbb{B}_{1}^{n}(r):=r\mathbb{B}_{1}^{n}$. We also define $\operatorname{rad}(\mathcal{K})=\sup_{\mathbf{u}\in\mathcal{K}}\|\mathbf{u}\|_{2}$. Recall that $\Sigma^{n,*}_{s}:=(\Sigma^{n}_{s})^{\mathbb{S}}$ is the set of all $s$-sparse signals in $\mathbb{S}^{n-1}$.
We refer to complex numbers with absolute value $1$ as the phase-only values. For a vector $\mathbf{u}=[u_{i}]\in\mathbb{C}^{n}$, we work with the $\ell_{p}$-norm $\|\mathbf{u}\|_{p}=(\sum_{i}|u_{i}|^{p})^{1/p}$ ($p\geq 1$), max norm $\|\mathbf{u}\|_{\infty}=\max_{i}|u_{i}|$, and zero “norm” $\|\mathbf{u}\|_{0}$ that counts the number of non-zero entries. Further given $\mathbf{v}=[v_{i}]\in\mathbb{C}^{n}$, we work with the inner product $\langle\mathbf{u},\mathbf{v}\rangle=\mathbf{u}^{*}\mathbf{v}=\sum_{i=1}^{n}\overline{u_{i}}v_{i}$ and the Hadamard product $\mathbf{u}\odot\mathbf{v}=(u_{1}v_{1},u_{2}v_{2},\cdots,u_{n}v_{n})^{\top}$. For a complex matrix $\mathbf{A}=\mathbf{B}+\mathbf{C}\mathbf{i}$ with $\mathbf{i}$ reserved for the imaginary unit, we will use $\mathbf{A}^{\Re}$ or $\Re(\mathbf{A})$ to denote its real part $\mathbf{B}$, and $\mathbf{A}^{\Im}$ or $\Im(\mathbf{A})$ to denote its imaginary part $\mathbf{C}$. For a random variable $X$ we define the sub-Gaussian norm as
$\|X\|_{\psi_{2}}=\inf\{t>0:\mathbbm{E}\exp(X^{2}/t^{2})<2\}$
and the sub-exponential norm as
$\|X\|_{\psi_{1}}=\inf\{t>0:\mathbbm{E}\exp(|X|/t)<2\}.$
For independent zero-mean sub-Gaussian variables $\{X_{i}\}_{i=1}^{N}$, there exists absolute constant $C$ such that
@@
The sub-Gaussian norm of a random vector $\mathbf{X}\in\mathbb{R}^{n}$ is defined as $\|\mathbf{X}\|_{\psi_{2}}=\sup_{\mathbf{v}\in\mathbb{S}^{n-1}}\|\mathbf{v}^{\top}\mathbf{X}\|_{\psi_{2}}$. We refer readers to [50, Sec. 2] for details on these definitions and properties.
[…]
We say $I_{1}$ is small enough (or sufficiently small) if $I_{1}\leq c_{1}$ for some suitably small constant $c_{1}$. Conversely, it is large enough (or sufficiently large) if $I_{1}\geq C_{1}$ for some suitably large constant $C_{1}$. We say $I_{1}$ is bounded away from 0 if $I_{1}\geq c_{1}$ for some $c_{1}>0$.
2.1 Linear Compressed Sensing
Let $\mathcal{U}$ be a cone in $\mathbb{R}^{n}$, we say $\mathbf{A}$ satisfies RIP over $\mathcal{U}$ with distortion $\delta>0$, denoted by $\mathbf{A}\sim\operatorname{RIP}(\mathcal{U},\delta)$, if